# A Cross-Scene Benchmark for Open-World Drone Active Tracking

### v1

@@ A.3 Reinforcement Learning in Visual Tracking
Reinforcement learning (RL) is commonly used in visual object tracking [69, 45, 71]. Song et al. [52] propose a decision-making mechanism based on hierarchical reinforcement learning (HRL), which achieves state-of-the-art performance while maintaining a balance between accuracy and computational efficiency. However, the actions generated by reinforcement learning in the aforementioned work cannot directly influence the camera’s viewpoint, thereby failing to fully leverage the decision-making capabilities. Real-world applications increasingly require robust tracking in highly dynamic scenes, motivating researchers to explore reinforcement learning agents for effectively synchronizing visual perception and decision-making in VAT tasks. Dionigi et al. [18] demonstrate the feasibility of reinforcement learning for drone VAT missions. However, the assumption of a fixed-forward perspective limits its applicability in real-world tasks.
@@ A.4 Curriculum Learning in Robot Control
Curriculum Learning (CL) is a training strategy that mimics a human curriculum by training models on simpler subsets of data at first and gradually expanding to larger and more difficult subsets of data until they are trained on the entire dataset. As for robot control, reinforcement learning training is difficult due to the complexity of the training scenarios and the large action spaces. Therefore, curriculum learning is often required to reduce the difficulty of agent training. For instance, many works improve the walking ability of legged robots by adjusting terrain parameters through curriculum learning [47, 35]. Other studies improve the pushing and grasping performance of robotic arms by progressively increasing task difficulty [32, 55, 39].
@@ Appendix B More Details of DAT Benchmark
Scenario Construction. In this study, three scenarios: citystreet, downtown, and lake are derived from real-world locations. The corresponding map segments for these scenarios are extracted from OpenStreetMap (OSM) [2], and road information, such as the number of lanes and intersection rules, is edited using the JOSM tool [1]. Then configurations are then converted into webots-compatible assets. Specifically, the citystreet scenario is based on a small town in Los Angeles, the downtown scenario is derived from Manhattan, and the lake scenario is modeled after Wolf Lake Memorial Park in Indiana. In contrast, the village, desert, and farmland maps possess complex and unique features that are not adequately captured by OpenStreetMap (OSM) data. For example, the village map features mountainous terrain with tunnels, while the farmland map is characterized by diverse multicolored patterns. To overcome these limitations, this study utilizes Creo software [26] to model detailed scene elements, which are then integrated into the webots for constructing realistic maps.
@@
Vision Camera Image Mat Images captured by the camera Default sensor
LiDAR LidarCloud vector2000 Point cloud from RPLIDAR A2 (m) Obstacle avoidance
@@
[…]
Gyroscope Angular vector3 Angular velocity (rad/s) Robot posture stabilization
IMU Angle vector3 Euler angles (rad) Robot posture stabilization
Orientation vector4 Quaternion representation Robot posture stabilization
Table 8: State parameters of DAT benchmark.
@@
cameraFov double camera field of view($rad$)
@@
cameraMidGlobalPos vector3d coordinates of the camera center mapped to the ground in the world coordinate system
carMidGlobalPos vector3d coordinates of the vehicle center in the world coordinate system
cameraMidPos vector3d coordinates of the camera center in the world coordinate system
carDronePosOri vector4d 1D orientation + 3D position of vehicle in the drone coordinate system
@@
Table 9: Reward parameters of DAT benchmark. The homogeneous transformation matrices (HTM) $\texttt{T}_{\texttt{ct}}$ and $\texttt{T}_{\texttt{ct}}$ are 4$\times$4 square matrices. Therefore, their data type $\text{double}[16]$ corresponds to a double array of length 16.
Targets. All tracking target illustrations are presented in Fig. 6. Specifically, Fig. 6(a) presents automobile and motorbike tracking targets, including passenger vehicles (the first seven cars), buses, trucks, trailers, and motorcycles (such as scooters and motorbikes). These two categories of tracking targets leverage Simulation of Urban Mobility (SUMO) [30] for road behavior modeling and interaction management with other targets. In contrast, Fig. 6(b)-(d) display pedestrian, wheeled robot, and legged robot tracking targets, respectively. These three types of targets utilize SUMO paths for position initialization and rely on specific controllers for action and behavior management.
Sensors. In the real world VAT tasks, a single camera cannot ensure the agent’s stability and robustness. Thus, integration with other sensors is often required. The DAT benchmark provides common sensors that can obtain the drone’s state parameters relative to the world coordinate system. The drone’s position and velocity are determined using GPS, while its acceleration is measured by an accelerometer, providing essential self-referential data for visual navigation tasks. Angular velocity is recorded via a gyroscope, and Euler angles obtained from the IMU are converted into quaternions to facilitate state estimation and ensure orientation stability. Additionally, the RPLIDAR A2, provided by DAT, generates point cloud data, which supports tasks such as obstacle avoidance and navigation by delivering detailed environmental information. The specific sensors, their parameters, and potential tasks are summarized in Tab. 8.
Additional Parameters. The training process of VAT agents often requires additional parameters for effective reward design. To facilitate this, DAT benchmark provides 4 categories comprising a total of 13 parameters, supporting diverse reward design strategies, as detailed in Tab. 9.
@@
Next is the homogeneous transformation matrix (HTM). In the reward design, coordinate transformations are often required to express physical quantities within a unified coordinate system, enabling consistent calculations. For example, prior studies [18, 33, 17] transform the position, velocity, and acceleration of targets into the tracker’s coordinate system to construct rewards. To support such operations, DAT benchmark provides $\texttt{T}_{\texttt{ct}}$, the HTM mapping the drone camera coordinate system to the world coordinate system, and $\texttt{T}_{\texttt{tw}}$, the HTM mapping the tracking target’s coordinate system to the world coordinate system.
@@
Lastly, for ease of model training in simulations, reward design often depends on some privileged information, i.e., variables that are almost impossible to obtain in real-world settings. Thus, DAT benchmark also provides such adaptations. For example, carMidGlobalPos gives the target’s position in the world coordinate system, and carDronePosOri represents the target’s orientation and position relative to the drone coordinate system, frequently used in VAT reward design [18, 33, 17]. Furthermore, information on the target’s movement direction and type is provided, enabling training the drone’s predictive ability for target movement.
Figure 7: Schemetic diagram of reward acquisition under fixed-forward perspective. (a) Illustration of the mapping between the image plane $LU_{1}LD_{1}RD_{1}RU_{1}$ and the target plane $LU_{G1}LD_{G1}RD_{G1}RU_{G1}$, along with the tracker body-fixed frame $\{t\}$, whose $x$-axis is aligned with the forward flight direction of the drone. (b) Schematic of the reward acquisition plane. As this plane is rectangular, the design of the distance-based reward becomes effective.
Task Configuration. We encapsulate the scenes, tasks, and data augmentation into Python classes, and provide 3 different environment classes for different algorithm requirements. The base environment class directly interacts with webots and is designed to support asynchronous reinforcement learning algorithms, such as the asynchronous advantage actor-critic (A3C) algorithm [36]. The Gymnasium environment class wraps the base environment class into a Gymnasium [54] interface, enabling direct compatibility with popular reinforcement learning libraries, such as Stable-Baselines3 [43] and Tianshou [61] for efficient algorithm development and evaluation. The parallel environment class encapsulates the base environment class to enable parallel execution, providing direct support for synchronous algorithms, such as proximal policy optimization (PPO) [49] and soft actor-critic (SAC) [23]. Additionally, the scenario selection, tracker and target configuration, SUMO parameters, task additional parameters, and randomization methods can all be efficiently customized through a JSON configuration file.
Appendix C More Details of Proposed R-VAT
@@ C.1 Theoretical Proof of Reward Design
Current reinforcement learning-based VAT methods assume a fixed forward-view perspective [73, 33, 17, 18] and design rewards based on physical distance. Distance-based reward design approaches fail to accurately reflect the performance of the agent in tracking tasks, especially in real-world applications that require a variable top-down perspective. In this section, we first demonstrate the effectiveness of distance-based reward design under the assumption of a fixed forward-view perspective. We then demonstrate that, under a top-down perspective, these approaches can result in scenarios where targets farther from the image center are assigned higher rewards than those closer to it.
The distance-based reward design approaches leverage the physical distance of the tracking target relative to the tracker to define the reward. While the exact reward formulations in existing methods differ, the underlying principle and the associated potential issues remain consistent. Therefore, we take the reward formulation from [18] as an example for the subsequent theoretical proof. The reward in [18] is expressed along the x, y, and z directions as follows:
$\small\begin{array}[]{rl}r_{x}(\cdot)\!&=\max{(0,1-\mid y_{x}(\cdot)-d_{r}\mid)},\\ r_{y}(\cdot)\!&=\max{(0,1-\mid\frac{2}{A_{FoV}}\arctan{(\frac{y_{y}(\cdot)}{y_{x}(\cdot)})}\mid)},\\ r_{z}(\cdot)\!&=\max{(0,1-\mid\frac{2}{A_{FoV}}\arctan{(\frac{y_{z}(\cdot)}{y_{x}(\cdot)})}\mid)},\end{array}$ (12)
where the reward is defined in the tracker body-fixed frame, as illustrated in Fig. 7(a).
Figure 8: Schematic diagram of pinhole imaging.
Effectiveness under a fixed forward-view perspective. As illustrated in Fig. 7(a), when the tracker adopts a fixed-forward perspective, the image plane is parallel to the plane containing the target. Based on the principle of pinhole imaging (see Fig. 8), the quadrilateral $LU_{1}LD_{1}RD_{1}RU_{1}$ on the image plane is geometrically similar to the quadrilateral $LU_{G1}LD_{G1}RD_{G1}RU_{G1}$ on the target plane, which forms a rectangle. Furthermore, the principle of pinhole imaging can also be utilized to establish the mapping range between the image plane and the target plane.
Figure 9: Schemetic diagram of reward acquisition under top-down perspective. (a) Illustration of the mapping between the image plane $LU_{2}LD_{2}RD_{2}RU_{2}$ and the target plane $LU_{G2}LD_{G2}RD_{G2}RU_{G2}$. (b) Failure case of the distance-based reward design in the $x$ direction and the tracker body-fixed frame $\{t\}$, which is consistent with the definition in [18]. When the target is located in the trapezoid $abcd$, a target farther from the image center may receive a higher reward. (c) Failure case in the $y$ direction. When the target is located in triangular regions $a_{1}b_{1}c_{1}$ and $a_{2}b_{2}c_{2}$, points farther from the center point may receive a higher reward.
To simplify the analysis of the reward function shown in Eq. 12, we define a function $f(x)=\!1-\!\!\mid\!A\arctan x\!\mid$ where $A\!=\!\frac{2}{A_{FoV}}$ and $x\!=\!\frac{y_{y}(\cdot)}{y_{x}(\cdot)}$. The function is monotonically decreasing when $x\!\!>\!\!0$ and monotonically increasing when $x\!\!<\!\!0$. Specifically, given a fixed $y_{x}$, the relationship holds:
$\forall y_{y1}\!,y_{y2}\!\in\![-\frac{W}{2},\frac{W}{2}]\text{, if }\mid\!\!y_{y1}\!\!\mid>\mid\!\!y_{y2}\!\!\mid\!\text{, }f(\!\frac{y_{y1}}{y_{x}}\!)\!\!<\!\!f(\!\frac{y_{y2}}{y_{x}}\!),$ (13)
$\forall y_{z1}\!,y_{z2}\!\in\![-\frac{H}{2},\frac{H}{2}]\text{, if }\mid\!\!y_{z1}\!\!\mid>\mid\!\!y_{z2}\!\!\mid\!\text{, }f(\!\frac{y_{z1}}{y_{x}}\!)\!\!<\!\!f(\!\frac{y_{z2}}{y_{x}}\!).$
Consequently, across the entire image, the reward decreases as the distance of the target from the center of the image increases. This demonstrates that Eq. 12 provides an effective measure of drone tracking performance in both the $y$ and $z$ directions under the assumption of a fixed forward-view perspective (See Fig. 7(b)).
Figure 10: The zoomed-in illustration of the triangular region $I_{OG}RU_{G2}RD_{G2}$ shown in Fig. 9(c). The point $P_{T}$ represents the coordinate point of the tracker’s projection onto the ground.
Figure 11: Schematic diagram of the optical frustum in a camera. (a) The camera’s optical frustum $O_{c}\!-\!CRU_{G2}LU_{G2}D$ and its intersection with the ground, forming the quadrilateral $LU_{G2}LD_{G2}RD_{G2}RU_{G2}$. This structure demonstrates the existence of the triangular region $a_{1}b_{1}c_{1}$ depicted in Fig. 9(c). (b) The zoomed-in illustration of the triangular regions $O_{c}EF$ and $O_{c}RU_{G2}LU_{G2}$, which are used in the theoretical proof presented in Eq. 17. (c) The zoomed-in illustration of the triangular regions $ORU_{G2}LU_{G2}$ and $NGH$, which are used in the theoretical proof presented in Eq. 18.
[…]
$\small S_{a^{\prime}bcd^{\prime}}\!=\!\{(x,y)\!\mid\!(x,y)\!\in\!\mathbb{R}^{2}\text{ and }(x,y)\text{ inside }a^{\prime}bcd^{\prime}\}.$ (14)
For any point $B\in S_{a^{\prime}bcd^{\prime}}$, it holds that $d(B)<d(A)$.
Moreover, according to Eq. 12, the reward $r_{x}(\cdot)$ in the $x$-direction is determined by $d_{x}(\cdot)=\mid y_{x}(\cdot)-d_{r}\mid$, which represents the distance in the x-direction of the target point from the image center. Based on this metric, the symmetric point $A^{\prime\prime}$ of point $A$ with respect to $I_{OG}$ and the iso-distance line segment $l_{ad}$ can be constructed. All points on $l_{ad}$ receive the same $x$-direction reward $r_{x}$ as point $A$.
Besides, the reward function $r_{x}$ in Eq. 12 decreases monotonically with increasing $d_{x}$. Thus, for any $d_{x}(B)>d_{x}(A)$, it follows that $r_{x}(B)<r_{x}(A)$. Accordingly, the set of points strictly inside trapezoid $abcd$ can be defined as:
$\small S_{abcd}\!=\!\{(x,y)\!\mid\!(x,y)\!\in\!\mathbb{R}^{2}\text{ and }(x,y)\text{ inside }abcd\}.$ (15)
For any $B\in S_{abcd}$, it follows that $r_{x}(B)<r_{x}(A)$, and $d(B)<d(A)$ as $B$ also lies within trapezoid $a^{\prime}bcd^{\prime}$. Therefore, it becomes evident that a point $A$, which is farther from the image center, may receive a higher reward than a point $B$, which is closer to the center. This highlights that the reward design in Eq. 12 fails to accurately represent tracking performance in the x-direction.
Failure Cases in the y-direction under Top-Down Perspective. In the $y$-direction, Fig. 9(c) illustrates another failure case of the reward, which occurs in the triangular regions $a_{1}b_{1}c_{1}$ and $a_{2}b_{2}c_{2}$. To ensure the validity of subsequent proofs, we first need to prove the existence of the triangular regions $a_{1}b_{1}c_{1}$ and $a_{2}b_{2}c_{2}$. This can be achieved by demonstrating that $\angle_{RU_{G2}OLU_{G2}}\!<\!FoV$ (See Sec. C.1 for theoretical proof), which ensures that the aforementioned triangular regions do not vanish due to parallel bounding edges. Subsequently, we analyze the failure case of the distance-based reward using the region $a_{1}b_{1}c_{1}$ as an example. A zoomed-in view is shown in Fig. 10.
As shown in Fig. 10, a target point $B$ is located within the region $a_{1}b_{1}c_{1}$, while another target point $A$ coincides with point $c_{1}$. According to the definition in Eq. 12, we can define a function $\theta(\cdot)=\arctan(\frac{y_{y}(\cdot)}{y_{x}(\cdot)})$, which represents the angle between the line segment $l_{P_{T}I_{OG}}$ and the line connecting the target coordinate point and the image center $I_{OG}$. For the target point $B$, $\theta(B)$ is defined as $\theta(B)=\angle_{BP_{T}I_{OG}}$, while for point $A$, $\theta(A)=\frac{FoV}{2}$. When the target point $B$ is within the region $a_{1}b_{1}c_{1}$, it holds that $\theta(B)<\theta(A)$. Furthermore, given that $\theta\!>\!0$, the function $r_{y}(\cdot)$ exhibits a monotonically decreasing behavior concerning $\theta$. As a result, it follows that $r_{y}(B)>r_{y}(A)$.
In addition, to determine the distance from point $B$ to the image center, a line segment $l_{BB^{\prime}}$ is constructed through point $B$ such that $l_{BB^{\prime}}\!\parallel\!l_{b_{1}A}$. Therefore, the inequality holds: $\frac{\mid l_{I_{OG}A}\mid}{\mid l_{I_{OG}RD_{G2}}\mid}\!<\!\frac{\mid l_{I_{OG}B^{\prime}}\mid}{\mid l_{I_{OG}RD_{G2}}\mid}$. Based on the definition of the distance from a point to the image center in Fig. 11, this implies that $d(B)>d(A)$. Consequently, the target point $B$, which is farther from the image center, receives a higher reward in the $y$-direction compared to the closer point $A$. This result demonstrates that the distance-based reward design fails to accurately reflect tracking performance in the $y$-direction.
Theoretical proof of angular relationships. To demonstrate that distance-based reward design is invalid in the top-down perspective (as discussed in Sec. C.1), it is first necessary to prove:
$\small\angle_{RU_{G2}OLU_{G2}}<FoV.$ (16)
The proof is based on the optical frustum shown in Fig. 11.
As shown in Fig. 11, the optical frustum $O_{c}\!-\!CRU_{G2}LU_{G2}D$ of the camera is depicted. The base quadrilateral $CRU_{G2}LU_{G2}D$ of the frustum is a rectangle. The line $l_{O_{c}O^{\prime}}$, which connects the apex $O_{c}$ to the ground center $O^{\prime}$, is perpendicular to the base plane. The intersection of the frustum with the ground plane produces the quadrilateral $LU_{G2}LD_{G2}RD_{G2}RU_{G2}$. This quadrilateral represents the projection of the image plane onto the ground plane.
Thus, under the aforementioned notation, the proof of Eq. 16 can be divided into the following two steps. First, we demonstrate that when the point $RD_{G2}$ coincides with the apex $O_{c}$, the base angle of triangle $ORU_{G2}LU_{G2}$ is greater than that of triangle $O_{c}EF$, as follows:
$\small\angle_{ORU_{G2}LU_{G2}}>\angle_{O_{c}EF}.$ (17)
Next, we establish that as $RD_{G2}$ moves from point $C$ to point $O_{c}$ along the line segment $l_{CO_{c}}$, the base angle of triangle $ORU_{G2}LU_{G2}$ decreases monotonically. Formally, this can be expressed as:
$\small\forall RD_{G21},RD_{G22}\in l_{CO_{c}}\text{, if }\mid l_{CRD_{G21}}\!\mid>\mid l_{CRD_{G22}}\!\mid\text{,}\\ \text{ then }\angle_{ORU_{G21}LU_{G21}}\!<\!\angle_{ORU_{G22}LU_{G22}}$ (18)
As shown in Fig. 11(b), it is given that $\mid\!l_{EO^{{}^{\prime}}}\!\!\mid=\mid\!l_{RU_{G2}H}\!\!\mid$ and $\mid\!\!l_{O_{c}H}\!\!\mid>\mid\!\!l_{O_{c}O^{{}^{\prime}}}\!\!\!\mid$. Based on these conditions, the following inequality is satisfied: $\frac{\mid l_{O_{c}O^{{}^{\prime}}}\mid}{\mid l_{EO^{{}^{\prime}}}\mid}\!\!<\!\!\frac{\mid l_{O_{c}H}\mid}{\mid l_{RU_{G2}H}\mid}$. Moreover, the base angle of triangle $O_{c}RU_{G2}LU_{G2}$ and $O_{c}EF$ can be expressed as:
$\angle_{O_{c}EF}$ $=\arctan{(\frac{\mid l_{O_{c}O^{{}^{\prime}}}\mid}{\mid l_{EO^{{}^{\prime}}}\mid})},$ (19)
$\angle_{O_{c}RU_{G2}LU_{G2}}$ $=\arctan{(\frac{\mid l_{O_{c}H}\mid}{\mid l_{RU_{G2}H}\mid})}.$
Given that the function $\arctan(\cdot)$ is monotonically increasing, it follows that: $\angle_{O_{c}RU_{G2}LU_{G2}}\!>\!\angle_{O_{c}EF}$.
As shown in Fig. 11(c), a line $l_{RD_{G2}R^{\prime}}$ is constructed through point $RD_{G2}$, perpendicular to $l_{RU_{G2}LU_{G2}}$. The length of this perpendicular segment is equal to the height of the trapezoid, i.e., $\mid l_{NH}\mid=\mid l_{RD_{G2}R^{{}^{\prime}}}\mid$. Consequently, the base angle of $ORU_{G2}LU_{G2}$ can be expressed as:
$\small\angle_{ORU_{G2}LU_{G2}}=\arctan{(\frac{\mid l_{NH}\mid}{\mid l_{RU_{G2}R^{{}^{\prime}}}\mid})}.$ (20)
In the trapezoid $LU_{G2}LD_{G2}RD_{G2}RU_{G2}$, the length of line segment $l_{RU_{G2}R^{{}^{\prime}}}$ is given by the following equation:
$\small\mid\!l_{RU_{G2}R^{{}^{\prime}}}\!\mid=\frac{1}{2}(\mid\!l_{RU_{G2}LU_{G2}}\!\mid\!-\!\mid\!l_{RD_{G2}LD_{G2}}\!\mid).$ (21)
Additionally, in triangle $O_{c}CD$, the relationship exists:
$\small\frac{\mid l_{RD_{G2}LD_{G2}}\mid}{\mid l_{CD}\mid}=1-\frac{\mid l_{NG}\mid}{\mid l_{O_{c}G}\mid}.$ (22)
By combining Eq. 21 and Eq. 22, and given that $\mid\!l_{RU_{G2}LU_{G2}}\!\mid=\mid\!l_{CD}\!\mid$ holds in rectangle $CRU_{G2}LU_{G2}D$, the length of line segment $l_{RU_{G2}R^{{}^{\prime}}}$ can be determined as:
$\small\mid l_{RU_{G2}R^{{}^{\prime}}}\mid=\frac{1}{2}\frac{\mid l_{NG}\mid\cdot\mid l_{CD}\mid}{\mid l_{O_{c}G}\mid}.$ (23)
In triangle $NGH$, by applying the law of cosines, the length of line segment $l_{NH}$ can be calculated as:
$\small\mid\!l_{NH}\!\mid=\sqrt{\mid\!l_{NG}\!\mid^{2}\!+\!\mid\!l_{GH}\!\mid^{2}\!-2\mid\!l_{NG}\!\mid\mid\!l_{GH}\!\mid\!\cos\theta}.$ (24)
By combining Eq. 23 and Eq. 24, the following relationship can be derived:
$\small\frac{\mid l_{NH}\mid}{\mid l_{RU_{G2}R^{{}^{\prime}}}\mid}=c\sqrt{1+x^{2}-2x\cos\theta},$ (25)
where $c=\frac{2\mid l_{O_{c}G}\mid}{\mid l_{CD}\mid}$ and $x=\frac{\mid l_{GH}\mid}{\mid l_{NG}\mid}$. Furthermore, in the right triangle $O_{c}O^{{}^{\prime}}G$, the value of $\cos\theta$ is given by: $\cos\theta=\frac{\mid l_{GH}\mid}{2\mid l_{O_{c}G}\mid}$. Since point $RD_{G2}$ moves along segment $l_{CO_{c}}$, it follows that $\mid l_{NG}\mid\in[0,\mid l_{O_{c}G}\mid]$, which implies:
$\small x>\cos\theta.$ (26)
We can further define function $g(x)=x^{2}\!-\!2x\cos\theta\!+\!1$ and its derivative $g^{{}^{\prime}}(x)=2(x-\cos\theta)$. According to Eq. 26, $g^{{}^{\prime}}(x)>0$, which indicates that $g(x)$ is monotonically increasing with respect to $x$.
Additionally, since $h(\cdot)=c\sqrt{\cdot}$ and $\arctan(\cdot)$ are monotonically increasing, and $x=\frac{\mid l_{GH}\mid}{\mid l_{NG}\mid}$ is monotonically decreasing with respect to $\mid l_{NG}\mid$, the monotonicity rule for composite functions ensures that $\angle_{ORU_{G2}LU_{G2}}$ in Eq. 20 is monotonically decreasing regarding $\mid l_{NG}\mid$. Therefore, the minimum value of $\angle_{ORU_{G2}LU_{G2}}$ is achieved when $\mid\!l_{NG}\!\mid$ reaches its maximum value, which equals $\mid\!l_{O_{c}G}\!\mid$, i.e., when $RD_{G2}$ coincides with $O_{c}$. The minimum value is given by:
$\small\min\{\angle_{ORU_{G2}LU_{G2}}\}=\angle_{O_{c}RU_{G2}LU_{G2}},$ (27)
which implies that: $\angle_{RU_{G2}OLU_{G2}}<FoV$.

### v2

@@ A.3 Reinforcement learning in Visual Tracking
Reinforcement learning (RL) is widely used in large language models hu2024dynamic and robot control leggedgym to improve exploration performance. It is also commonly applied in visual object tracking actiondecision ; drlis ; drlvideo . Song et al. Song2020OnlineDB propose a decision-making mechanism based on hierarchical reinforcement learning (HRL), which achieves state-of-the-art performance while maintaining a balance between accuracy and computational efficiency. However, the actions generated by reinforcement learning in the aforementioned work cannot directly influence the camera’s viewpoint, thereby failing to fully leverage the decision-making capabilities. Real-world applications increasingly require robust tracking in highly dynamic scenes, motivating researchers to explore reinforcement learning agents for effectively synchronizing visual perception and decision-making in VAT tasks. Dionigi et al. dvat demonstrate the feasibility of reinforcement learning for drone VAT missions. However, the assumption of a fixed-forward perspective limits its applicability in real-world tasks.
@@ A.4 Curriculum Learning in Robot Control
Curriculum Learning (CL) is a training strategy that mimics a human curriculum by training models on simpler subsets of data at first and gradually expanding to larger and more difficult subsets of data until they are trained on the entire dataset. CL is widely used in large language models DBLP:conf/ijcai/0004HHLZCHWY0XZ25 and robot control. As for robot control, reinforcement learning training is difficult due to the complexity of the training scenarios and the large action spaces. Therefore, curriculum learning is often required to reduce the difficulty of agent training. For instance, many works improve the walking ability of legged robots by adjusting terrain parameters through curriculum learning leggedgym ; scirobot_distill . Other studies improve the pushing and grasping performance of robotic arms by progressively increasing task difficulty luo2020accelerating ; vaehrens2022learning ; or2023curriculum .
@@ Appendix B More Details of DAT Benchmark
More details of the digital twin tool. Our digital twin tool is based on the osm_importer tool in the webots simulation software. Users first need to download the map description file (.osm file) for a specific area from the OpenStreetMap website. Then, the tool preprocesses the map according to the configuration, modifying information such as the number of lanes and lane directions, and converts the processed file into a road network file (.net.xml file) that can be read by SUMO. Following this, the tool adds traffic lights and intersection traffic rules to the road network based on the configuration, ensuring that the traffic flow operates correctly when the map is converted into a 3D scene. Finally, the tool reads the road, vegetation, and building information and converts them into PROTO assets for webots, which can then be correctly recognized and used by the DAT benchmark.
Scenario Construction. Among the DAT scenes, three scenarios: citystreet, downtown, and lake are directly derived from real-world locations with the digital twins tool. Specifically, the citystreet scenario is based on a small town in Los Angeles, the downtown scenario is derived from Manhattan, and the lake scenario is modeled after Wolf Lake Memorial Park in Indiana. In contrast, the village, desert, and farmland maps possess complex and unique features that are not adequately captured by OpenStreetMap (OSM) data. For example, the village map features mountainous terrain with tunnels, while the farmland map is characterized by diverse multicolored patterns. To overcome these limitations, we use Creo software creo to model detailed scene elements, which are then integrated into the webots for constructing realistic maps.
Table 8: State parameters of DAT benchmark.
@@
Vision Camera Image Mat Images Default sensor
LiDAR LidarCloud vector2000 Point cloud (m) Obstacle avoidance
@@
Accelerometer Acc vector3 Acceleration ($\text{m/s}^{2}$) Visual navigation
Gyroscope Angular vector3 Angular velocity (rad/s) Posture stabilization
IMU Angle vector3 Euler angles (rad) Posture stabilization
Orientation vector4 Quaternion representation Posture stabilization
Table 9: Reward parameters of DAT benchmark. The homogeneous transformation matrices (HTM) $\texttt{T}_{\texttt{cw}}$ and $\texttt{T}_{\texttt{tw}}$ are $\times4$ square matrices. Therefore, their data type $\text{double}[16]$ corresponds to a double array of length 16.
@@
cameraFov double camera field of view(radrad)
@@
cameraMidGlobalPos vector3d ground projection of camera center mapped in the world frame
carMidGlobalPos vector3d coordinates of the vehicle center in the world frame
cameraMidPos vector3d coordinates of the camera center in the world frame
carDronePosOri vector4d 1D orientation + 3D position of vehicle in the drone frame
@@
Targets. All tracking target illustrations are presented in Fig. 9. Specifically, Fig. 9(a) presents automobile and motorbike tracking targets, including passenger vehicles (the first seven cars), buses, trucks, trailers, and motorcycles (such as scooters and motorbikes). These two categories of tracking targets leverage Simulation of Urban Mobility (SUMO) sumo for road behavior modeling and interaction management with other targets. In contrast, Fig. 9(b)-(d) display pedestrian, wheeled robot, and legged robot tracking targets, respectively. These three types of targets utilize SUMO paths for position initialization and rely on specific controllers for action and behavior management.
Figure 10: Diagram for the theoretical proof of Proposition 1.
Sensors. In the real world VAT tasks, a single camera cannot ensure the agent’s stability and robustness. Thus, integration with other sensors is often required. The DAT benchmark provides common sensors that can obtain the drone’s state parameters relative to the world coordinate system. The drone’s position and velocity are determined using GPS, while its acceleration is measured by an accelerometer, providing essential self-referential data for visual navigation tasks. Angular velocity is recorded via a gyroscope, and Euler angles obtained from the IMU are converted into quaternions to facilitate state estimation and ensure orientation stability. Additionally, the RPLIDAR A2, provided by DAT, generates point cloud data, which supports tasks such as obstacle avoidance and navigation by delivering detailed environmental information. The specific sensors, parameters and potential tasks are in Table 8.
Additional Parameters. The training process of VAT agents often requires additional parameters for effective reward design. To facilitate this, DAT benchmark provides 4 categories comprising a total of 13 parameters, supporting diverse reward design strategies, as detailed in Table 9.
@@
Next is the homogeneous transformation matrix (HTM). In the reward design, coordinate transformations are often required to express physical quantities within a unified coordinate system, enabling consistent calculations. For example, prior studies dvat ; e2e ; cvat transform the position, velocity, and acceleration of targets into the tracker’s coordinate system to construct rewards. To support such operations, DAT benchmark provides $\texttt{T}_{\texttt{cw}}$, the HTM mapping the drone camera coordinate system to the world coordinate system, and $\texttt{T}_{\texttt{tw}}$, the HTM mapping the tracking target’s coordinate system to the world coordinate system.
@@
Lastly, for ease of model training in simulations, reward design often depends on some privileged information, i.e., variables that are almost impossible to obtain in real-world settings. Thus, DAT benchmark also provides such adaptations. For example, carMidGlobalPos gives the target’s position in the world coordinate system, and carDronePosOri represents the target’s orientation and position relative to the drone coordinate system, frequently used in VAT reward design dvat ; e2e ; cvat . Furthermore, information on the target’s direction and type is provided.
Task Configuration. We encapsulate the scenes, tasks, and domain randomization into Python classes, and provide 3 different environment classes for different algorithm requirements. The base environment class directly interacts with webots and is designed to support asynchronous reinforcement learning algorithms, such as the asynchronous advantage actor-critic (A3C) algorithm a3c . The Gymnasium environment class wraps the base environment class into a Gymnasium gymnasium interface, enabling direct compatibility with popular reinforcement learning libraries, such as Stable-Baselines3 stable_bl3 and Tianshou tianshou for efficient algorithm development and evaluation. The parallel environment class encapsulates the base environment class to enable parallel execution, providing direct support for synchronous algorithms, such as proximal policy optimization (PPO) ppo and soft actor-critic (SAC) sac . Additionally, the scenario selection, tracker and target configuration, SUMO parameters, task additional parameters, and randomization methods can all be efficiently customized through a JSON configuration file.
Appendix C More Details of Proposed GC-VAT
@@ C.1 Theoretical Proof of Reward Design
Theoretical proof of Proposition 1. Consider two points $A$ and $B$ on the image symmetry axis (in Fig. 10), which are symmetric with respect to the image center $O$. The projections of these points onto the ground are denoted as $A^{\prime}$, $B^{\prime}$ and $O^{\prime}$, respectively. Take a point $C^{\prime}$ on line segment $O^{\prime}B^{\prime}$ such that the Euclidean distance $d(O^{\prime},C^{\prime})=d(A^{\prime},O^{\prime})$.
Given:
1. In the image plane, the deviation $\phi(\cdot,\cdot)$ of point $A$ and $B$ from the image center $O$ is the same, i.e. $\phi(A,O)=\phi(B,O)$.
2. In the projection plane, the Euclidean distance from $A^{\prime}$ and $C^{\prime}$ to the ideal position $O^{\prime}$ are equal, i.e. $d(A^{\prime},O^{\prime})=d(O^{\prime},C^{\prime})$.
It is evident that:
1. For any point $D^{\prime}$ on line segment $B^{\prime}C^{\prime}$, the following relationship holds: $d(O^{\prime},D^{\prime})>d(A^{\prime},O^{\prime})$.
2. The corresponding point $D$ in the image lies on line segment $BC$, and thus $\phi(D,O)<\phi(A,O)$.
Thus, it is clear that the actual distance between the target and the ideal position is inconsistent with the deviation of the target from the image center in the image.
Theoretical proof of Remark 1. According to Proposition 1, the following relationship between the Euclidean distance and the deviation holds:
$\small\exists P_{1},P_{2}\in\mathcal{I}_{p}\text{, s.t. }\phi_{1}<\phi_{2}\text{, }d_{1}>d_{2},$ (8)
where $\phi_{i}=\phi(P_{i},C_{g})$ and $d_{i}=d(P_{i},C_{g})$. Therefore, for a distance-based reward function $\mathcal{R}_{d}(\cdot)$ that satisfies the Reward Design Principle, it follows that:
$\exists P_{1},P_{2}\in\mathcal{I}_{p}\text{, s.t. }\phi_{1}<\phi_{2}\text{, }\mathcal{R}_{d}(d_{1})<\mathcal{R}_{d}(d_{2}).$ (9)
@@
Scene citystreet desert village downtown lake farmland
Steps (M) 19.2 13.4 21.3 19.8 9.9 9.2
Table 11: Transition steps across different scenes.
Scene citystreet desert village downtown lake farmland
T (M) 10.0 6.2 8.0 10.3 5.6 4.1